# Personalized Federated Learning with Heat-Kernel Enhanced Tensorized Multi-View Clustering

### v1

@@ 4.1.6 Tensorized Local Clustering Algorithm
Algorithm 2 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local)
@@
Input: Local tensor data $\mathcal{X}_{i}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{P}_{g},\mathbf{Q}_{g},\mathbf{R}_{g}\}$, personalization parameters $\{\lambda_{i},\rho_{i}\}$
@@
$\mathcal{G}_{i}^{(0)}\leftarrow\lambda_{i}\mathcal{G}_{g}+(1-\lambda_{i})\mathcal{G}_{i}^{local}$
$\mathbf{P}_{i}^{(0)},\mathbf{Q}_{i}^{(0)},\mathbf{R}_{i}^{(0)}\leftarrow$ similarly personalized initialization
$\mathbf{v}_{i}^{(0)}\leftarrow\frac{1}{s_{i}}\mathbf{1}$
@@
  Calculate $\delta_{j\ell h}^{(i)}$ using Eq. 61
@@
  $\mathcal{A}_{i}\leftarrow\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i}$
@@
  Update $\mathcal{G}_{i}^{(t)}$ using Eq. 68
  Update $\mathbf{P}_{i}^{(t)}$ using Eq. 69
  Update $\mathbf{Q}_{i}^{(t)}$ using Eq. 70
  Update $\mathbf{R}_{i}^{(t)}$ using Eq. 71
@@
  Compute $J_{i}^{tensor}$ using Eq. 63
@@
return $\mathcal{U}_{i}^{(t)},\mathcal{G}_{i}^{(t)},\mathbf{P}_{i}^{(t)},\mathbf{Q}_{i}^{(t)},\mathbf{R}_{i}^{(t)},\mathbf{v}_{i}^{(t)}$
@@ 4.2 Federated Aggregation with Personalization
The federated learning framework coordinates local heat-kernel enhanced clustering results while preserving client personalization. The aggregation process operates on cluster centers and view importance weights.
@@ 4.2.1 Global Objective with Personalization
$J_{global}=\sum_{i=1}^{M}\left[\omega_{i}J_{i}^{local}(U_{i},A_{i},V_{i})+\gamma\mathcal{R}_{person}(A_{i},A_{g})+\eta\mathcal{R}_{view}(V_{i},V_{g})\right]$ (73)
@@
  • $\omega_{i}$ represents client-specific weights based on data quality and quantity
  • $\mathcal{R}_{person}(A_{i},A_{g})=\sum_{h,k}\|A_{i}^{h}-A_{g}^{h}\|_{F}^{2}$ ensures coordination between local and global centers
  • $\mathcal{R}_{view}(V_{i},V_{g})=\|V_{i}-V_{g}\|_{2}^{2}$ maintains view weight consistency
@@ 4.2.2 Privacy-Preserving Aggregation
Each client $i$ computes and shares:
$S_{i}^{centers}$ $=\left\{\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}x_{j}^{h,(i)},\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\right\}_{k,h}$ (74)
$S_{i}^{views}$ $=\left\{\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{KED}_{i}^{h}(x_{j}^{h,(i)},a_{k}^{h,(i)})\right\}_{h}$ (75)
$S_{i}^{quality}$ $=J_{i}^{local}\text{ and }n_{i}$ (76)
@@
$A_{g}^{h,(t+1)}=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{centers,h}}{\sum_{i=1}^{M}\omega_{i}^{(t)}\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}}$ (77)
$V_{g}^{(t+1)}=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}V_{i}^{(t)}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (78)
@@
$\omega_{i}^{(t)}=\frac{\exp(-\tau J_{i}^{local,(t)})}{\sum_{j=1}^{M}\exp(-\tau J_{j}^{local,(t)})}$ (79)
@@ 4.2.3 Personalized Update Mechanism
$A_{i}^{h,(t+1)}$ $=\lambda_{i}A_{i}^{h,(t)}+(1-\lambda_{i})A_{g}^{h,(t+1)}$ (80)
$V_{i}^{(t+1)}$ $=\rho_{i}V_{i}^{(t)}+(1-\rho_{i})V_{g}^{(t+1)}$ (81)
where $\lambda_{i},\rho_{i}\in[0,1]$ are client-specific personalization parameters that can be learned adaptively:
$\lambda_{i}^{(t+1)}=\lambda_{i}^{(t)}-\beta\nabla_{\lambda_{i}}J_{i}^{local}(\lambda_{i}A_{i}^{(t)}+(1-\lambda_{i})A_{g}^{(t)})$ (82)
@@ 4.3.1 Tensorized Global Objective with Personalization
$J_{tensor\_global}=\sum_{i=1}^{M}\left[\omega_{i}J_{i}^{tensor}(\mathcal{U}_{i},\mathcal{G}_{i},\mathbf{P}_{i},\mathbf{Q}_{i},\mathbf{R}_{i},\mathbf{v}_{i})+\gamma\mathcal{R}_{tensor}(\mathcal{G}_{i},\mathbf{P}_{i},\mathbf{Q}_{i},\mathbf{R}_{i})+\eta\mathcal{R}_{view}(\mathbf{v}_{i},\mathbf{v}_{g})\right]$ (83)
@@
$\mathcal{R}_{tensor}$ $=\|\mathcal{G}_{i}-\mathcal{G}_{g}\|_{F}^{2}+\|\mathbf{P}_{i}-\mathbf{P}_{g}\|_{F}^{2}+\|\mathbf{Q}_{i}-\mathbf{Q}_{g}\|_{F}^{2}+\|\mathbf{R}_{i}-\mathbf{R}_{g}\|_{F}^{2}$ (84)
$\mathcal{R}_{view}(\mathbf{v}_{i},\mathbf{v}_{g})$ $=\|\mathbf{v}_{i}-\mathbf{v}_{g}\|_{2}^{2}$ (85)
@@ 4.3.2 Privacy-Preserving Tensor Aggregation
Each client $i$ computes and shares:
$S_{i}^{core}$ $=\left\{\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\mathcal{G}_{i,k,:,:}\right\}$ (86)
$S_{i}^{factors}$ $=\left\{\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\mathbf{P}_{i,k,:},\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\mathbf{Q}_{i,:,\ell},\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\mathbf{R}_{i,:,h}\right\}$ (87)
$S_{i}^{views}$ $=\left\{\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{TKED}_{i}(j,k,h)\right\}_{h}$ (88)
$S_{i}^{quality}$ $=J_{i}^{tensor}\text{ and }n_{i}$ (89)
@@
$\mathcal{G}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{core}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (90)
$\mathbf{P}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{factors,P}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (91)
$\mathbf{Q}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{factors,Q}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (92)
$\mathbf{R}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{factors,R}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (93)
$\mathbf{v}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}\mathbf{v}_{i}^{(t)}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (94)
@@ 4.3.3 Tensorized Personalized Update Mechanism
$\mathcal{G}_{i}^{(t+1)}$ $=\lambda_{i}^{core}\mathcal{G}_{i}^{(t)}+(1-\lambda_{i}^{core})\mathcal{G}_{g}^{(t+1)}$ (95)
$\mathbf{P}_{i}^{(t+1)}$ $=\lambda_{i}^{P}\mathbf{P}_{i}^{(t)}+(1-\lambda_{i}^{P})\mathbf{P}_{g}^{(t+1)}$ (96)
$\mathbf{Q}_{i}^{(t+1)}$ $=\lambda_{i}^{Q}\mathbf{Q}_{i}^{(t)}+(1-\lambda_{i}^{Q})\mathbf{Q}_{g}^{(t+1)}$ (97)
$\mathbf{R}_{i}^{(t+1)}$ $=\lambda_{i}^{R}\mathbf{R}_{i}^{(t)}+(1-\lambda_{i}^{R})\mathbf{R}_{g}^{(t+1)}$ (98)
$\mathbf{v}_{i}^{(t+1)}$ $=\rho_{i}\mathbf{v}_{i}^{(t)}+(1-\rho_{i})\mathbf{v}_{g}^{(t+1)}$ (99)
where $\lambda_{i}^{core},\lambda_{i}^{P},\lambda_{i}^{Q},\lambda_{i}^{R},\rho_{i}\in[0,1]$ are component-specific personalization parameters.
4.4 Complete Tensorized Federated Algorithm
Algorithm 3 Personalized Federated Heat-Kernel Enhanced Tensorized Multi-View Clustering
@@ 4.4 Complete Tensorized Federated Algorithm
 Initialize global tensor factors $\mathcal{G}_{g}^{(0)},\mathbf{P}_{g}^{(0)},\mathbf{Q}_{g}^{(0)},\mathbf{R}_{g}^{(0)}$ and view weights $\mathbf{v}_{g}^{(0)}$
@@
  Broadcast: Send $\mathcal{G}_{g}^{(t)},\mathbf{P}_{g}^{(t)},\mathbf{Q}_{g}^{(t)},\mathbf{R}_{g}^{(t)},\mathbf{v}_{g}^{(t)}$ to all clients
@@
  for each client $i\in\{1,\ldots,M\}$ do
@@
   Organize $X_{i}=\{X_{i}^{h}\}_{h=1}^{s_{i}}$ into tensor $\mathcal{X}_{i}$
@@
   $\mathcal{G}_{i}^{(0)}\leftarrow\lambda_{i}^{core}\mathcal{G}_{g}^{(t)}+(1-\lambda_{i}^{core})\mathcal{G}_{i}^{local}$
   $\mathbf{P}_{i}^{(0)}\leftarrow\lambda_{i}^{P}\mathbf{P}_{g}^{(t)}+(1-\lambda_{i}^{P})\mathbf{P}_{i}^{local}$
   $\mathbf{Q}_{i}^{(0)}\leftarrow\lambda_{i}^{Q}\mathbf{Q}_{g}^{(t)}+(1-\lambda_{i}^{Q})\mathbf{Q}_{i}^{local}$
   $\mathbf{R}_{i}^{(0)}\leftarrow\lambda_{i}^{R}\mathbf{R}_{g}^{(t)}+(1-\lambda_{i}^{R})\mathbf{R}_{i}^{local}$
   $\mathbf{v}_{i}^{(0)}\leftarrow\rho_{i}\mathbf{v}_{g}^{(t)}+(1-\rho_{i})\mathbf{v}_{i}^{local}$
@@
   for $e=1,2,\ldots,E$ do
    Compute $\delta_{j\ell h}^{(i)}$ using Eq. 61
    Reconstruct $\mathcal{A}_{i}\leftarrow\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i}$
    Update $\mathcal{U}_{i}$ using Eq. 67
    Update $\mathcal{G}_{i}$ using Eq. 68
    Update $\mathbf{P}_{i}$ using Eq. 69
    Update $\mathbf{Q}_{i}$ using Eq. 70
    Update $\mathbf{R}_{i}$ using Eq. 71
    Update $\mathbf{v}_{i}$ using Eq. 72
@@
   Compute $S_{i}^{core},S_{i}^{factors},S_{i}^{views},S_{i}^{quality}$
@@
  Compute client weights $\omega_{i}$ using tensor quality measures
  Update $\mathcal{G}_{g}^{(t+1)}$ using Eq. 90
  Update $\mathbf{P}_{g}^{(t+1)}$ using Eq. 91
  Update $\mathbf{Q}_{g}^{(t+1)}$ using Eq. 92
  Update $\mathbf{R}_{g}^{(t+1)}$ using Eq. 93
  Update $\mathbf{v}_{g}^{(t+1)}$ using Eq. 94
@@
  Update tensor-specific personalization parameters $\{\lambda_{i}^{core},\lambda_{i}^{P},\lambda_{i}^{Q},\lambda_{i}^{R},\rho_{i}\}$

### v2

@@ 6.4 Local Clustering Algorithms
Algorithm 2 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local FedHK-PARAFAC2)
@@
Input: Local tensor data $\mathcal{X}_{[\ell]}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{P}_{g},\mathbf{Q}_{g},\mathbf{R}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, $g$ is the global model index
@@
$\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{[\ell]}\mathcal{G}_{g}+(1-\lambda_{[\ell]})\mathcal{G}_{[\ell]}$
$\mathbf{P}_{[\ell]}^{(0)},\mathbf{Q}_{[\ell]}^{(0)},\mathbf{R}_{[\ell]}^{(0)}\leftarrow$ similarly personalized initialization
$v_{[\ell]}^{(0)}\leftarrow\frac{1}{s(\ell)}\mathbf{1}$
@@
  Calculate $\delta_{[\ell]ij}^{h}$ using Eq. 61
@@
  $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$
@@
  Update $\mathcal{G}_{[\ell]}^{(t)}$ using Eq. 69
  Update $\mathbf{P}_{[\ell]}^{(t)}$ using Eq. 72
  Update $\mathbf{Q}_{[\ell]}^{(t)}$ using Eq. 73
  Update $\mathbf{R}_{[\ell]}^{(t)}$ using Eq. 74
@@
  Compute $J_{[\ell]}^{tensor}$ using Eq. 64
@@
return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{P}_{[\ell]}^{(t)},\mathbf{Q}_{[\ell]}^{(t)},\mathbf{R}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$
Algorithm 3 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local FedHK-Tucker)
Input: Local tensor data $\mathcal{X}_{[\ell]}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{U}_{g},\mathbf{V}_{g},\mathbf{W}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, $g$ is the global model index
Initialize:
$\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{[\ell]}\mathcal{G}_{g}+(1-\lambda_{[\ell]})\mathcal{G}_{[\ell]}$
$\mathbf{U}_{[\ell]}^{(0)},\mathbf{V}_{[\ell]}^{(0)},\mathbf{W}_{[\ell]}^{(0)}\leftarrow$ similarly personalized initialization
$v_{[\ell]}^{(0)}\leftarrow\frac{1}{s(\ell)}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute tensorized heat-kernel coefficients
  Calculate $\delta_{[\ell]ij}^{h}$ using Eq. 61
  // Reconstruct cluster centers from decomposition
  $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$
  // Update memberships
  Update $U_{[\ell]}^{*^{(t)}}$ using Eq. 79
  // Update tensor factors using ALS
  Update $\mathcal{G}_{[\ell]}^{(t)}$ using Eq. 80
  Update $\mathbf{U}_{[\ell]}^{(t)}$ using Eq. 83
  Update $\mathbf{V}_{[\ell]}^{(t)}$ using Eq. 84
  Update $\mathbf{W}_{[\ell]}^{(t)}$ using Eq. 85
  // Update view weights
  Update $v_{[\ell]}^{(t)}$ using Eq. 81
  // Check convergence
  Compute $J_{[\ell]}^{tensor}$ using Eq. 75
until convergence or max iterations
return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{U}_{[\ell]}^{(t)},\mathbf{V}_{[\ell]}^{(t)},\mathbf{W}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$
@@ 6.5 Federated Aggregation with Personalization
Personalization in federated learning allows clients to maintain individualized models while benefiting from shared knowledge. In the context of heat-kernel enhanced multi-view clustering, personalization enables clients to adapt global clustering insights to their unique data distributions. To enable personalization, we introduce client-specific parameters that modulate the influence of global model updates on local clustering results.
@@ 6.5.1 Global Objective with Personalization
This section extends the federated aggregation framework to incorporate personalization for the heat-kernel enhanced multi-view clustering. First, we define the global objective function that balances local clustering quality with personalization constraints. The global model maintains shared cluster centers and view weights, while clients adapt these based on their local data characteristics.
The global model parameters are defined as:
  • Global cluster centers: $A_{\text{global}}=\{a_{kj}^{h,(\text{global})}\}_{k=1,j=1,h=1}^{c,d_{h},s}$
  • Global view weights: $V_{\text{global}}=\{v_{h}^{(\text{global})}\}_{h=1}^{s}$
The global federated objective with personalization is formulated as follows.
Definition 1 (Federated Heat-Kernel Enhanced Multi-View Clustering with Personalization).
@@
$J_{\text{global}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{FedHK}^{(\ell)}(U_{[\ell]},A_{[\ell]},V_{[\ell]})+\gamma\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}}\right]$ (86)
@@
  • $\omega_{\ell}$ represents client-specific weights based on data quality and quantity
  • $\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}}=\sum_{h,k}\|A_{[\ell]}^{h}-A_{\text{global}}^{h}\|_{F}^{2}$ ensures coordination between local and global centers
  • $\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$ maintains view weight consistency
@@
We call this framework Federated Heat-Kernel Enhanced Multi-View Clustering with Personalization (FedHK-MVC-Person). Simply recall the local objective from FedHK-PARAFAC2 in Eq. 64 and FedHK-Tucker in Eq. 75 to instantiate $J_{FedHK}^{(\ell)}$ accordingly.
@@ 6.5.2 Privacy-Preserving Aggregation
Each client $\ell$ computes and shares:
$S_{[\ell]}^{centers}$ $=\left\{\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}x_{[\ell]ij}^{h},\sum_{j=1}^{n_{i}}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\right\}_{kj,h}$ (87)
$S_{[\ell]}^{views}$ $=\left\{\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c}(\mu_{[\ell]ik}^{*})^{m}\text{KED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})\right\}_{h}$ (88)
$S_{[\ell]}^{quality}$ $=J_{[\ell]}\text{ and }n(\ell)$ (89)
@@
$A_{\text{global}}^{h,(t+1)}=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{\ell}^{centers,h}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}\sum_{j=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}}$ (90)
$V_{\text{global}}^{(t+1)}=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)V_{\ell}^{(t)}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (91)
@@
$\omega_{\ell}^{(t)}=\frac{\exp(-\tau J_{[\ell]}^{(t)})}{\sum_{\ell=1}^{M}\exp(-\tau J_{[\ell]}^{(t)})}$ (92)
@@ 6.5.3 Personalized Update Mechanism
$A_{\ell}^{h,(t+1)}$ $=\lambda_{\ell}A_{\ell}^{h,(t)}+(1-\lambda_{\ell})A_{\text{global}}^{h,(t+1)}$ (93)
$V_{\ell}^{(t+1)}$ $=\rho_{\ell}V_{\ell}^{(t)}+(1-\rho_{\ell})V_{\text{global}}^{(t+1)}$ (94)
where $\lambda_{\ell},\rho_{\ell}\in[0,1]$ are client-specific personalization parameters that can be learned adaptively:
$\lambda_{\ell}^{(t+1)}=\lambda_{\ell}^{(t)}-\beta\nabla_{\lambda_{\ell}}J_{[\ell]}(\lambda_{\ell}A_{\ell}^{(t)}+(1-\lambda_{\ell})A_{\text{global}}^{(t)})$ (95)
@@ 6.6.1 Tensorized Global Objective with Personalization
$J_{\text{tensor(global)}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{[\ell]}^{tensor}(\mathcal{U}_{[\ell]},\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]},\mathbf{V}_{[\ell]})+\gamma\mathcal{R}_{tensor}(\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})\right]$ (96)
@@
$\mathcal{R}_{tensor}$ $=\|\mathcal{G}_{[\ell]}-\mathcal{G}_{\text{global}}\|_{F}^{2}+\|\mathbf{P}_{[\ell]}-\mathbf{P}_{\text{global}}\|_{F}^{2}+\|\mathbf{Q}_{[\ell]}-\mathbf{Q}_{\text{global}}\|_{F}^{2}+\|\mathbf{R}_{[\ell]}-\mathbf{R}_{\text{global}}\|_{F}^{2}$ (97)
$\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})$ $=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$ (98)
@@ 6.6.2 Privacy-Preserving Tensor Aggregation
Each client $\ell$ computes and shares:
$S_{[\ell]}^{core}$ $=\left\{\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c}(\mu_{[\ell]ik}^{*})^{m}\mathcal{G}_{[\ell]i,k,:,:}\right\}$ (99)
$S_{[\ell]}^{factors}$ $=\left\{\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\mathbf{P}_{[\ell]i,j,:},\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\mathbf{Q}_{[\ell]i,:,h},\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ij}^{*})^{m}\mathbf{R}_{[\ell]i,:,h}\right\}$ (100)
$S_{[\ell]}^{views}$ $=\left\{\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}_{[\ell]}(i,k,h)\right\}_{h}$ (101)
$S_{[\ell]}^{quality}$ $=J_{[\ell]}^{tensor}\text{ and }n(\ell)$ (102)
@@
$\mathcal{G}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{core}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (103)
$\mathbf{P}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{factors,P}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (104)
$\mathbf{Q}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{factors,Q}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (105)
$\mathbf{R}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{factors,R}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (106)
@@ 6.6.3 Tensorized Personalized Update Mechanism
$\mathcal{G}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{core}\mathcal{G}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{core})\mathcal{G}_{\text{global}}^{(t+1)}$ (107)
$\mathbf{P}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{P}\mathbf{P}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{P})\mathbf{P}_{\text{global}}^{(t+1)}$ (108)
$\mathbf{Q}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{Q}\mathbf{Q}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{Q})\mathbf{Q}_{\text{global}}^{(t+1)}$ (109)
$\mathbf{R}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{R}\mathbf{R}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{R})\mathbf{R}_{\text{global}}^{(t+1)}$ (110)
where $\lambda_{\ell}^{core},\lambda_{\ell}^{P},\lambda_{\ell}^{Q},\lambda_{\ell}^{R},\rho_{\ell}\in[0,1]$ are component-specific personalization parameters.
Algorithm 4 provides a comprehensive framework for implementing personalized FedHK-PARAFAC2. In order to achieve a personalized FedHK-Tucker, two changes must be made. Initially, adjustments are made to the tensor factor implication stage and its localization tensorized HKC. Secondly, modifications are made to the tensor aggregation stage through the implementation of Tucker decomposition.
Algorithm 4 Personalized FedHK-PARAFAC2
@@
 Initialize global tensor factors $\mathcal{G}_{\text{global}}^{(0)},\mathbf{P}_{\text{global}}^{(0)},\mathbf{Q}_{\text{global}}^{(0)},\mathbf{R}_{\text{global}}^{(0)}$ and view weights $V_{\text{global}}^{(0)}$
@@
  Broadcast: Send $\mathcal{G}_{\text{global}}^{(t)},\mathbf{P}_{\text{global}}^{(t)},\mathbf{Q}_{\text{global}}^{(t)},\mathbf{R}_{\text{global}}^{(t)},V_{\text{global}}^{(t)}$ to all clients
@@
  for each client $\ell\in\{1,\ldots,M\}$ do
@@
   Organize $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ into tensor $\mathcal{X}_{[\ell]}$
@@
   $\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{core}\mathcal{G}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{core})\mathcal{G}_{[\ell]}$
   $\mathbf{P}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{P}\mathbf{P}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{P})\mathbf{P}_{[\ell]}$
   $\mathbf{Q}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{Q}\mathbf{Q}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{Q})\mathbf{Q}_{\ell}$
   $\mathbf{R}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{R}\mathbf{R}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{R})\mathbf{R}_{[\ell]}$
   $A_{[\ell]}^{(0)}\leftarrow\rho_{\ell}A_{\text{global}}^{(t)}+(1-\rho_{\ell})A_{[\ell]}$
   $V_{[\ell]}^{(0)}\leftarrow\rho_{\ell}V_{\text{global}}^{(t)}+(1-\rho_{\ell})V_{[\ell]}$
@@
   for $e=1,2,\ldots,E+1$ do
    Compute $\delta_{[\ell]ij}^{h}$ using Eq. 61
    Reconstruct $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$
    Perform Algorithm 2
@@
   Compute $S_{[\ell]}^{core},S_{[\ell]}^{factors},S_{[\ell]}^{views},S_{[\ell]}^{quality}$
@@
  Compute client weights $\omega_{\ell}$ using tensor quality measures
  Update $\mathcal{G}_{\text{global}}^{(t+1)}$ using Eq. 103
  Update $\mathbf{P}_{\text{global}}^{(t+1)}$ using Eq. 104
  Update $\mathbf{Q}_{\text{global}}^{(t+1)}$ using Eq. 105
  Update $\mathbf{R}_{\text{global}}^{(t+1)}$ using Eq. 106
  Update $A_{\text{global}}^{(t+1)}$ using Eq. 90
  Update $V_{\text{global}}^{(t+1)}$ using Eq. 91
@@
  Update tensor-specific personalization parameters $\{\lambda_{\ell}^{core},\lambda_{\ell}^{P},\lambda_{\ell}^{Q},\lambda_{\ell}^{R},\rho_{\ell}\}$